OpenStudy (anonymous):

some one help with this integral. i will write it out:

hero (hero):

Somebody please play the Jeopardy song

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

\[\int\limits_{}^{} \sin x dx\]

OpenStudy (anonymous):

\[- \cos x\]

OpenStudy (anonymous):

What a tease :(

OpenStudy (anonymous):

-cos x + C